OpenStudy (anonymous):

Can someone help me find a solution to this integral? (x^3) / sqrt((x^2) + 1)

OpenStudy (anonymous):

the whole denominator is under the radical

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{x^3}{\sqrt{x^2+1}}dx\]Ehhh you could take several different approaches to solve this one. Trigonometric Substitution. U-sub. Integration by parts. What do you prefer?

OpenStudy (anonymous):

u-sub please:)

zepdrix (zepdrix):

Umm ok I think we can make that work. Let's borrow a couple powers of x from the numerator like this:\[\Large\rm \int\limits x^2\frac{x~dx}{\sqrt{x^2+1}}\] We'll make the substitution: \(\Large\rm u=\sqrt{x^2+1}\) Then taking a derivative gives us:\[\Large\rm du=\frac{x~dx}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

can you show me more steps I can't quite understand how you got the final answer...

OpenStudy (anonymous):

ah nvm I got it

zepdrix (zepdrix):

\[\Large\rm \int\limits\limits x^2\frac{x~dx}{\sqrt{x^2+1}}=\int\limits x^2 du\]That takes care of a lot of the stuff for us. More steps? Sure. I kinda rushed through that.

OpenStudy (anonymous):

learning along :D

zepdrix (zepdrix):

Where we having trouble? How I got from u to du? Or why I chose that for u in the first place?

OpenStudy (anonymous):

so all I need to do is solve for the integral of X^2 and get the integral plus c?

zepdrix (zepdrix):

No no no. We can't integrate x^2 with our differential du. We have to also replace the x^2 with something in u.

OpenStudy (anonymous):

@zepdrix how do you know what u to pick to get the answer the least painful way?

zepdrix (zepdrix):

\[\Large\rm u=\sqrt{x^2+1}\]Square each side,\[\Large\rm u^2=x^2+1\]Subtract 1 from each side and voila!\[\Large\rm u^2-1=x^2\]

zepdrix (zepdrix):

That takes some practice, just being able to eyeball it. I had initially thought that \(\Large\rm u=x^2+1\) would be a good substitution. But it leaves us with \(\Large\rm (stuff)^{3/2}\) in the numerator which would be difficult to deal with.

OpenStudy (xapproachesinfinity):

@study100 because when you do du it will give you that expression which is what we needed

zepdrix (zepdrix):

I think Trig Sub is the best approach if you're scratching your head. It probably ends up being a longer process. But it's a safe way to get through the problem if you can't `see` the u substitution. Sometimes they're hidden, like this one.

OpenStudy (anonymous):

@zepdrix ohhh I see :) so that's why you leave the radical with u . Thank you! @xapproachesinfinity Yep, but without knowing that in the first place, I guess I'd go with x^2 +1 and then find my way slowly to rad (x^2+1) , but it takes skills to get it in one go like zepdrix

zepdrix (zepdrix):

Well I did it the wrong way on paper as my first approach lol :) Took me a minute or two.

OpenStudy (xapproachesinfinity):

the first thing I thought about is rad(x^2+1)=u which is what he did here. but I made a mistake in differentiating that radical and got different result than he did. thus, I could see that x^2 separation to x.xdx

zepdrix (zepdrix):

Where you at Johnny boy? Confused still?

OpenStudy (anonymous):

what if it was +1 instead of -1 under the radical

OpenStudy (anonymous):

sorry -1

zepdrix (zepdrix):

It would work out largely the same. At least our du would be the same:\[\Large\rm \int\limits\limits x^2\frac{x~dx}{\sqrt{x^2+1}}=\int\limits\limits x^2~du\]Just the x^2 in front would change a lil bit.

zepdrix (zepdrix):

\[\Large\rm \int\limits\limits\limits x^2\frac{x~dx}{\sqrt{x^2-1}}=\int\limits\limits\limits x^2~du\]Woops*

OpenStudy (anonymous):

alrighty thanks for your help

OpenStudy (anonymous):

It's a bit different for me then. Everyone has a different thought process and I need to pratctice some more with u substitution. I guess you're naturally good with this so that's really great for you :)

OpenStudy (anonymous):

or you have lots of practice etc. Practice makes perfect.

OpenStudy (xapproachesinfinity):

well sometimes it works sometimes not and i got stuck as all people.

OpenStudy (xapproachesinfinity):

I actually i have gotten rusty in this stuff, i have been away from math. I'm recently trying to restore what i know LOL.

OpenStudy (xapproachesinfinity):

But of course practice is the key to this. it becomes second nature when you do it often lol

OpenStudy (anonymous):

Oh you're still really great now! :) That's amazing.

OpenStudy (anonymous):

Yes, I'm told that too lol :)

OpenStudy (xapproachesinfinity):

can you believe i'm starting over from precalc

OpenStudy (xapproachesinfinity):

Not really great, I make severe mistake in differentiation haha

OpenStudy (anonymous):

oh? you're good at integration too, which is pretty important in Calc 1 and 2

OpenStudy (xapproachesinfinity):

Next semester i will get back to Calc1. my favorite integral are those that involve trigs

OpenStudy (anonymous):

You're majoring in math? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

Yes! and you

OpenStudy (anonymous):

I'm just a calc student lol :) I'm going to be in college soon, bio major taking bio, prolly organic chem and calc 2

OpenStudy (xapproachesinfinity):

Oh! I see. Bio is an interesting major as well. Best of luck to you. we meet again! I got a class now. My name is Hassan by the way

OpenStudy (anonymous):

Nice meeting you, Hassan :) I'm Alice. Have fun in class!